5.6. What is a lower bound for the price of a six-month call option on a non-dividend-paying stock when the stock price is $80, the strike price is $75, and the risk-free interest rate is 10% per annum?
Answer #1

The lower bound will be

Stock price - present value of exercise price

80 - 75 / e ^ ( 10% * 6/12) = 80 - 71.34 = $8.66